# The Neglected Tails of Vision-Language Models

### 23 Jan 2024 v1

@@ F Generalization Performance of REAL
Table 14: REAL-Prompt generalizes across prompt templates. We show that REAL-Prompt (using the most frequent synonyms) can improve upon both OpenAI prompt templates [33] and LLM-generated templates such as DCLIP [26] and CuPL [32].
@@ G More Failures of Multimodal Systems
In Figure 8 and 9, we show more failure cases of state-of-the-art multimodal systems on tailed concepts identified by our frequency estimation method. These tailed concepts are randomly sampled from nine benchmark datasets and span across a variety of domains, including birds, flowers, fungi, snakes, frogs, fish, household items, and more. We qualitatively test the visual recognition abilities of two most popular visual chatbots: GPT4-V [48] (trained on proprietary data) and LLaVA1.5 [21] (trained on open-source data using a frozen CLIP image encoder). We also test the image generation abilities of two most popular generative models: DALL-E 3 [39] (trained on proprietary data) and Stable Diffusion XL [2] (trained on open-source data using a frozen CLIP text encoder). We observe that these systems fail to recognize or generate more than half of the tailed concepts we sampled. In particular, LLaVA1.5 and Stable Diffusion XL fail on all these tailed concepts, suggesting a large performance gap between proprietary and open-source multimodal systems.
Figure 8: State-of-the-art multimodal systems fail to recognize or generate tailed concepts (part 1). We show more failure cases of popular multimodal systems (GPT4-V [48], LLaVA1.5 [21], DALL-E 3 [39], and Stable Diffusion XL [2]) on tailed concepts sampled from standard benchmark datasets such as ImageNet [9], Flowers [28], Aircraft [24], and etc. For GPT4-V and LLaVA1.5, we include example images of incorrectly predicted classes to show that visual chatbots often misclassify rare concepts as some similar-looking yet more common concepts. We include a definition for each tailed concept to show that DALL-E 3 and Stable Diffusion (SD-XL) can fail to capture the correct colors, shapes, and other characteristics of these concepts.
Figure 9: State-of-the-art multimodal systems fail to recognize or generate tailed concepts (part 2). We show more failure cases of popular multimodal systems (GPT4-V [48], LLaVA1.5 [21], DALL-E 3 [39], and Stable Diffusion XL [2]) on tailed concepts sampled from standard benchmark datasets such as ImageNet [9], Flowers [28], Aircraft [24], and etc. For GPT4-V and LLaVA1.5, we include example images of incorrectly predicted classes to show that visual chatbots often misclassify rare concepts as some similar-looking yet more common concepts. We include a definition for each tailed concept to show that DALL-E 3 and Stable Diffusion (SD-XL) can fail to capture the correct colors, shapes, and other characteristics of these concepts.

### 02 Feb 2024 v2

@@ F Generalization Performance of REAL
Table 14: REAL-Prompt generalizes across prompt templates. We show that REAL-Prompt (using the most frequent synonyms) can improve upon both OpenAI prompt templates [35] and LLM-enriched templates such as DCLIP [28] and CuPL [34].
@@ G More Failures of Multimodal Systems
In Figure 8 and 9, we show more failure cases of state-of-the-art multimodal systems on tailed concepts identified by our frequency estimation method. These tailed concepts are randomly sampled from nine benchmark datasets and span across a variety of domains, including birds, flowers, fungi, snakes, frogs, fish, household items, and more. We qualitatively test the visual recognition abilities of two most popular visual chatbots: GPT-4V [50] (trained on proprietary data) and LLaVA1.5 [23] (trained on open-source data using a frozen CLIP image encoder). We also test the image generation abilities of two most popular generative models: DALL-E 3 [41] (trained on proprietary data) and Stable Diffusion XL [2] (trained on open-source data using a frozen CLIP text encoder). We observe that these systems fail to recognize or generate more than half of the tailed concepts we sampled. In particular, LLaVA1.5 and Stable Diffusion XL fail on all these tailed concepts, suggesting a large performance gap between proprietary and open-source multimodal systems.
Figure 8: State-of-the-art multimodal systems fail to recognize or generate tailed concepts (part 1). We show more failure cases of popular multimodal systems (GPT-4V [50], LLaVA1.5 [23], DALL-E 3 [41], and Stable Diffusion XL [2]) on tailed concepts sampled from standard benchmark datasets such as ImageNet [10], Flowers [30], Aircraft [26], and etc. For GPT-4V and LLaVA1.5, we include example images of incorrectly predicted classes to show that visual chatbots often misclassify rare concepts as some similar-looking yet more common concepts. We include a definition for each tailed concept to show that DALL-E 3 and Stable Diffusion (SD-XL) can fail to capture the correct colors, shapes, and other characteristics of these concepts.
Figure 9: State-of-the-art multimodal systems fail to recognize or generate tailed concepts (part 2). We show more failure cases of popular multimodal systems (GPT-4V [50], LLaVA1.5 [23], DALL-E 3 [41], and Stable Diffusion XL [2]) on tailed concepts sampled from standard benchmark datasets such as ImageNet [10], Flowers [30], Aircraft [26], and etc. For GPT-4V and LLaVA1.5, we include example images of incorrectly predicted classes to show that visual chatbots often misclassify rare concepts as some similar-looking yet more common concepts. We include a definition for each tailed concept to show that DALL-E 3 and Stable Diffusion (SD-XL) can fail to capture the correct colors, shapes, and other characteristics of these concepts.